OpenStudy (anonymous):

HELPPP how do you find 3 points for this equation 3x+2y=2?

OpenStudy (anonymous):

pick an \(x\) solve for \(y\) pick anything, but try to make it easy by picking numbers like 0 and 2 and 4

OpenStudy (anonymous):

3x+2y=2 2y=-3x+2 y=(-3x+2)/2 just plug in X and you will get Y I recommend you use a table

OpenStudy (anonymous):

@brooke6635 3x+2y=2 2y=-3x+2 y=(-3x+2)/2 just plug in X and you will get Y I recommend you use a table

OpenStudy (anonymous):

quyss i need a answer.......im checking my friends question if shes right @jk_16

OpenStudy (anonymous):

FINEEEEEEEEEEEEEEEEEEEEEEEEE!!!!!!!! >.<

OpenStudy (anonymous):

Just use a table of values example: y=(-3x+2)/2 x=0, y=-3*0 +2 / 2 2/2=1 (0,1)